OpenStudy (anonymous):

Find the solution of the differential equation dP/dt = sqrt(P*t) that satisfies the initial condition P(1)=5. Answer: P(t) = ....

OpenStudy (anonymous):

\[\Large \frac{ dP }{ dt } = \sqrt{Pt}\]

OpenStudy (anonymous):

\[\large \frac{ (dP)^2 }{ (dt)^2 } = Pt\]

OpenStudy (anonymous):

\[\large \sqrt{\frac{ 1 }{ P }}dP = \sqrt{t}dt\]

OpenStudy (anonymous):

integrate both sides and use initial conditions \[\large 2(5)^{\frac{1}{2}} = \frac{2}{3}(1)^{\frac{3}{2}} + C\]

OpenStudy (anonymous):

\[\large C = 2\sqrt{5} - \frac{2}{3}\]

OpenStudy (anonymous):

So, \[\large P(t) = {(\frac{2}{6}t^{\frac{2}{3}}+\sqrt{5} - \frac{2}{6})}^{2}\]

OpenStudy (anonymous):

Answer is wrong...please help me find my mistake

OpenStudy (anonymous):

@estudier

OpenStudy (anonymous):

\[\frac{dP}{dt}=\sqrt{Pt} \rightarrow \frac{dP}{\sqrt{P}}=\sqrt{t}dt\]

OpenStudy (anonymous):

rest like you did

OpenStudy (anonymous):

I agree with your integral, so maybe it is the algebra, still checking......

OpenStudy (anonymous):

@myko : that is what i did, just left out the step that included doing that

OpenStudy (anonymous):

after integral i get\[\large 2P^{\frac{1}{2}} = \frac{2}{3}t^{\frac{3}{2}} + C\]

OpenStudy (anonymous):

that's same what i got, now just put initial condition

OpenStudy (anonymous):

which i sub in t=1 and P=5 and i get: 2sqrt(5) - 2/3 = C

OpenStudy (anonymous):

do it like this: \[5=(\frac{1}{3}+C)^2\]

OpenStudy (anonymous):

@myko : we didnt get the same...i have a 2/3 in front of my t and you have 1/3

OpenStudy (anonymous):

and solve for C

OpenStudy (anonymous):

i devided by 2

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

\[5=\frac{1}{9}+\frac{2C}{3}+C^2\]

OpenStudy (anonymous):

so does C have two answers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[C= - \frac{1}{3}-\sqrt{5} \] \[C=\sqrt{5} - \frac{1}{3}\]

OpenStudy (anonymous):

i dont think it can have two answers as I have to type this in on an online submission thing and it only accepts one answer. if you look at my equation (13 min ago) starting with 2P^(1/2) = ... if you sub the initial conditions in you only get one answer for C

OpenStudy (anonymous):

but you are forgetting that your P is under sqrt

OpenStudy (anonymous):

Sorry, i know I'm probably being quite dumb but why does that affect it?

OpenStudy (anonymous):

maybe you could write it this way: \[P=(\frac{1}{3}t^{\frac{3}{2}}-\frac{1}{3} \pm \sqrt{5})^2\]

OpenStudy (anonymous):

the root thing is: P=(something)^2 so this something could be +something or -something. Since it is squared, you get same P

OpenStudy (anonymous):

i had just been typing it in wrong as t^(2/3) instead of t^{3/2}. That was the only mistake. NOOOOOOOOOOOOO! @myko thanks for all the help. greatly appreciated :)

OpenStudy (anonymous):

:). Glad you got it